OpenStudy (anonymous):

-2x-5=10 Solve the equation. Be sure to show all work.

OpenStudy (karatechopper):

add five to both equations..

OpenStudy (karatechopper):

i mean sides sorry..

OpenStudy (karatechopper):

-2x-5=10 -2x=15 x=-7.5